The first number in an arthimetic sequence is 20. The fifth number is 37. Find the second number.
Answer by checkley79(3341)

20,w,x,y,z,37
w+x+y+z=37-20
w+x+y+z=17
17/4=4.25 the average increase in numbers.
20+4.25=24.25 ans for the second number.